OpenStudy (anonymous):

an altitude h of a triangle is increased by a length m.how much must be taken from the corresponding base b so that the area of the new triangle is half of the origignal triangle

OpenStudy (valpey):

\[\frac{1}{2}b*h = 2*\frac{1}{2}(b-x)*(h+m)\]

OpenStudy (anonymous):

can u tell me the answer only please

OpenStudy (valpey):

\[x = \frac{-1}{2}\frac{b*h}{h+m} - b\]

OpenStudy (valpey):

Or maybe I meant: \[x = \frac{-1}{2}\frac{b*h}{h+m} + b\]

OpenStudy (anonymous):

answer? + b or - b

OpenStudy (anonymous):

I yhink the answer is : \[\frac{b(2m+h)}{2(h+m)}\]